# Non-perturbative determination of the sphaleron rate for first-order phase transitions

### v1

@@ Appendix B Semi-classical calculation in terms of Higgs condensate in perturbation theory
In Ref. Li et al. (2025) a semi-classical calculation151515 In this approach, a sphaleron field configuration is found at classical level within the 3D EFT. This corresponds to a leading order approximation in perturbation theory, which would receive corrections from a fluctuation determinant computed for the sphaleron field configuration. of the sphaleron rate has been formulated for $x\gg 0.1$ as well as for $x<0.1$. In the latter case
@@
At higher orders within the 3D EFT perturbation theory Farakos et al. (1994b), it has been shown that this condensate can be computed with remarkable agreement with the lattice result, for strong phase transitions at small $x$ Ekstedt et al. (2022, 2024).
@@
where the overall factor is simply the expectation value of the tree-level potential, and the remainder is a fit with coefficients $A=26.12$, $B=-2.145$, $C=0.4237$ and $D=0.00717$. This concludes our detour to perturbation theory.

### v2

@@ Appendix B Semi-classical calculation in terms of Higgs condensate in perturbation theory
In Ref. Li et al. (2025) a semi-classical calculation [^17] In this approach, a sphaleron field configuration is found at classical level within the 3D EFT. This corresponds to a leading order approximation in perturbation theory, which would receive corrections from a fluctuation determinant computed for the sphaleron field configuration. of the sphaleron rate has been formulated for $x\gg 0.1$ as well as for $x<0.1$. In the latter case
@@
At higher orders within the 3D EFT perturbation theory Farakos et al. (1994a), it has been shown that this condensate can be computed with remarkable agreement with the lattice result, for strong phase transitions at small $x$ Ekstedt et al. (2022, 2024).
@@
where the overall factor is simply the expectation value of the tree-level potential, and the remainder is a fit with coefficients $A=26.12$, B=2.145B=-2.145, $C=0.4237$ and $D=0.00717$. This concludes our detour to perturbation theory.